# EliGen: Entity-Level Controlled Image Generation with Regional Attention

### v2

@@ Abstract
Recent advancements in diffusion models have significantly advanced text-to-image generation, yet global text prompts alone remain insufficient for achieving fine-grained control over individual entities within an image. To address this limitation, we present EliGen, a novel framework for Entity-Level controlled Image Generation. We introduce regional attention, a mechanism for diffusion transformers that requires no additional parameters, seamlessly integrating entity prompts and arbitrary-shaped spatial masks. By contributing a high-quality dataset with fine-grained spatial and semantic entity-level annotations, we train EliGen to achieve robust and accurate entity-level manipulation, surpassing existing methods in both spatial precision and image quality. Additionally, we propose an inpainting fusion pipeline, extending EliGen’s capabilities to multi-entity image inpainting tasks. We further demonstrate EliGen’s flexibility by integrating it with other open-source models such as IP-Adapter, In-Context LoRA and MLLM, unlocking new creative possibilities. The source code, model, and dataset are published at the Project Page.
@@ 1 Introduction
Recently, diffusion models Rombach et al. (2022a); Esser et al. (2024); Labs (2024) have significantly advanced text-to-image generation, enabling the creation of high-quality images through textual descriptions. However, a single prompt is insufficient for precise image design control, as natural language cannot describe image content with pixel-level precision, and image models often struggle to accurately model spatial relationships and object quantities Podell et al. (2024). This pitfall significantly limits their broader applicability. Consequently, integrating diverse conditions into text-to-image models has become a central research focus. For example, recent studies like ControlNet Zhang et al. (2023) and IP-Adapter Ye et al. (2023) have introduced conditions such as depth maps, normal maps and reference images to control the layout and style of generated images. Nevertheless, these conditions operate at a global image level, rather than targeting specific entities within the image. As a result, they lack the capability to provide fine-grained and entity-level control. This limitation impedes the generation of complex, highly customized images, motivating us to investigate precise controlling approaches for entity-level control.
@@ 2 Related Work
Text-to-Image Diffusion Models Ho et al. (2020); Song et al. (2021); Li et al. (2024) have catalyzed the evolution of generative AI by generating high-quality images from textual descriptions. Latent Diffusion Models (LDMs) Rombach et al. (2022a); Podell et al. (2024) have emerged as a pivotal innovation, operating in a compressed latent space rather than directly in pixel space, which significantly reduces computational costs. These models primarily incorporate text conditions through the cross attention module within the UNet Ronneberger et al. (2015) architecture. The Diffusion Transformer (DiT) Peebles and Xie (2022) represents the first diffusion model to utilize a Transformer, integrating conditions into image latents via adaptive LayerNorm. SD3 Esser et al. (2024) and FLUX Labs (2024) are the state-of-the-art models that adopt the DiT, enhancing image generation by concatenating text and latent image modalities, thereby fully leveraging the semantic information of text to produce images of higher quality and finer detail. Despite the impressive capabilities of these models, generating images solely from text remains constrained. Recent studies have been exploring methods to introduce additional conditions Zhang et al. (2023); Mou et al. (2023); Ye et al. (2023), with entity-level control being one such sophisticated condition.
@@
To equip foundational models with fine-grained control capabilities, researchers originally explored train-free methods such as diffusion fusion Bar-Tal et al. (2023) and cross attention guidance Ma et al. (2024). However, these approaches often produce unstable results and encounter entity omission issues. To tackle these challenges, researchers introduced additional modules to encode spatial information for each entity using bounding boxes and conducted task-specific training. GLIGEN Li et al. (2023) is the first entity control model to undergo training, integrating an additional gated attention module and encoding bounding boxes through Fourier embedding. Subsequently, MIGC Zhou et al. (2024) introduces the shading aggregation controller and maps spatial information onto attention maps to guide cross attention processes. InstanceDiffusion Wang et al. (2024b) further categorizes spatial information from coarse to fine and designs unique tokenizers for each category. While these studies focus on designing complex structures to encode spatial information, this paper shifts the focus to designing and training regional attention mechanism without extra parameters and incorporating spatial information by controlling the regions where entity semantics exert influence.
@@ 3.1 Preliminaries
Entity-level controlled image generation aims at generating images based on both global and entity-level conditions. The diffusion model should ensure not only the accuracy of attributes and positions of each entity but also the overall harmony and quality of the image. The conditional inputs include a global text prompt $p$ and $L$ entity-level conditions. Each entity is composed of a local text prompt $p_{i}$ and spatial localization $l_{i}$, denoted as:
@@ 3.2 Regional Attention
Here, we construct the binary mask $\mathrm{M}$ to specify the regions that each token should attend to during attention, as shown in Figure 3c, which primarily includes four categories of regions. For consistency, we treat the global prompt as a global “entity”, with its entity mask being 1 for the entire image.
@@ 3.3 Dataset with Entity Annotation
For the entity control task, each training sample should consist of an image, a global prompt, and multiple entity annotations. Regarding the image source, prior works typically use open-source datasets like Laion LAION (2022) for images, but we observe that Laion’s image distribution differs significantly images generated by base model, leading to stylistic inconsistencies. Using Laion for training would distort the model distribution and degrade image quality (see the Supplementary Material for ablations), as the model would need to fit this mismatch. Therefore, we choose to directly use images generated by our baseline model, FLUX.1-dev, to avoid such distribution discrepancies.
The dataset generation pipeline and statistics are illustrated in the Supplementary Material. Starting with DiffusionDB, we randomly select a caption and use FLUX.1-dev to generate the source image. Then, we employ Multimodal Large Language Mode (MLLM) Wang et al. (2024a) to re-caption the image, forming the global prompt. Unlike previous methods relying on Grounding DINO Liu et al. (2023), we directly use Qwen2-VL 72B, which possesses the most powerful grounding capabilities across all MLLMs, to randomly select entities in the image and annotate their local prompts and localization (via rectangular masks). This ensures consistency between local and global prompts. Since our regional attention employs soft-coded spatial localization, training with rectangular masks is sufficient to enable the model to generalize to arbitrary-shaped masks.
@@ 4.1 Entity-Level Controlled Generation
In this scenario, entities are localized using bounding boxes. Our model converts bounding boxes to masks for input, while others utilize bounding box coordinates. The results are shown in Figure 4, where our model demonstrates superior performance in image quality, and entity accuracy. In the first two rows, simple tasks with distinct entities are handled by most methods except Multi-Diffusion. The last two rows highlight complex layouts where entities are intricately coupled. While other models struggle to maintain accuracy and coherence, our model consistently delivers high-quality, coherent images.
@@ 5.1 Experimental Setup
Following previous methods, we perform quantitative evaluations on the filtered COCO Lin et al. (2014); Zheng et al. (2023) 2017 validation set, comprising 2830 prompt-image pairs. Previous GLIGEN Li et al. (2023) utilized the mean average precision for evaluation, a metric that computes the area under the precision-recall curve. However, we observe its critical limitation: it treats background entities in generated images as negative samples, resulting in substantial inaccuracies. As a result, we decided not to use this metric. Instead, to thoroughly evaluate the models’ capabilities on the COCO benchmark, we assess their performance across the following dimensions.
@@ 5.2 Comparison on COCO Benchmark
The evaluation results of the models on the COCO benchmark are presented in Tables 1. For Entity Success Rate, we assessed all methods using Qwen2-VLWang et al. (2024a), InternVL2Chen et al. (2024), and CogVLM2 Hong et al. (2024). The results demonstrate that EliGen consistently outperformed other models, highlighting its robustness and high quality. In terms of mIoU, our model performs comparably to Instance-Diffusion and substantially better than other methods. The slightly lower mIoU of EliGen compared to Instance-Diffusion is primarily due to its soft-coding scheme for positional information. This scheme allows EliGen to automatically expand certain regions to ensure image coherence, as further illustrated in the Supplementary Material. In terms of both CLIP Score and Aesthetic Score, EliGen significantly outperforms other models, demonstrating substantial superiority in image quality.
@@ 6 Conclusion
To achieve image generation with entity-level controls, we propose regional attention into the DiT architecture to effectively integrate entity prompts and their locations. Furthermore, we construct a dataset with fine-grained entity information and finetune the regional attention mechanism, developing the EliGen model. EliGen demonstrates high generalization in entity-level controlled image generation, delivering superior image quality and entity accuracy compared to existing models. We further propose an inpainting fusion pipeline, extending EliGen’s functionality to image inpainting tasks. Moreover, by integrating EliGen with open-source models, we demonstrate its potential for creative applications, broadening its scope and versatility.
@@ Appendix B Inpainting Fusion Algorithm
Figure 11: Dialogue-based image design and editing with MLLM and our EliGen. Users can utilize MLLM to design the layout of the image they wish to generate or input the image they want to edit into MLLM. After MLLM generates the caption and entity information of the image, EliGen can produce the target image.
@@ Appendix C Potential Creative Extension
Figure 12: Training qualitative comparison between our constructed dataset and Laion dataset. The image pair is generated from the same noise, and the global prompt is “A person standing by the river.” The comparative results show that the images generated by the model trained on our proposed dataset significantly outperform those from the Laion dataset in quality.
@@ C.2 Entity Transfer with In-Context LoRA
In-Context LoRA is capable of performing few-shot image composition tasks conditioned on either global prompts or images, and can be applied in various scenarios such as portrait photography, product design, font design, and visual identity design. Our EliGen model further enhances these scenarios by incorporating entity-level control conditions, enabling more personalized image generation. Taking visual identity design as an example, we have successfully accomplished the entity transfer task by introducing EliGen, as illustrated in Figure 10. Given the entity defined by the source image, we can transfer it to any desired location in the target image through EliGen’s positional control capability, thereby achieving the intended visual identity.
@@ C.3 Dialogue-based Image Design and Editing with MLLM
The MLLM is capable of interpreting user-provided images and textual content to engage in dialogue. As a result, we can achieve dialogue-based image design and editing by leveraging the integration of MLLM and EliGen. Specifically, grounding-enabled MLLMs such as Qwen2-VL can be utilized to determine entity locations based on textual descriptions, enabling the generation of images solely through text using EliGen. Furthermore, the MLLM can be employed to comprehend images and design entity information in accordance with specific image editing requirements, thereby facilitating precise image editing in conjunction with EliGen. The qualitative results are shown in Figure 11.
@@ D.1 Training Qualitative Comparison for Dataset Image Source
The primary objective of training the EliGen model is to activate its capability for layout control, which necessitates the construction of a dataset annotated with entities. Moreover, the image distribution within the dataset should closely align with the output distribution of the model itself. If the distribution gap is substantial, a significant portion of the training loss would be attributed to this discrepancy. This would markedly impede the model’s convergence and degrade the quality of the generated images. Consequently, we opt to utilize the Flux-generated images for dataset creation, rather than directly employing images from the open-source Laion dataset LAION [2022].
We conduct training experiments using both our custom-created dataset and the Laion dataset, where the annotations for the Laion dataset a re s derived from the same pipeline described in Section A.1. As illustrated in Figure 12, under identical prompts and initial noise conditions, the model trained on our dataset demonstrates significantly superior image quality compared to the model trained on the Laion dataset. Compared to Flux-generated images, the latter exhibit significant visual clutter, inferior detail quality, and a notable deviation from the expected distribution.
@@ E.2 Continuously Varying Entity Positions
In Figure 15, we evaluated the control efficacy of EliGen using identical input noise and the same entity. The position of the “person” entity shifts progressively from left to right, with EliGen maintaining consistent control throughout the process. To preserve the identity of the person during generation, we integrate the IP-Adapter Team [2024] method. This demonstrates EliGen’s potential to generate coherent image sequences, such as those involving continuous movement of entities through positional control or complex actions orchestrated via coordinated control of multiple entities.
@@ E.3 Unreasonable Positional Inputs
incorrect spatial relationships between entities refer to scenarios where the relative positions of entities are inaccurate, such as being too far apart or too close together. We evaluated two cases of this inputs, as depicted in Figure 16. In both examples, the correct positioning of the “person” entity necessitates an overlap with another entity to accurately convey the action being performed. For example, in the prompt “A person is playing tennis,” the “person” should ideally be holding a “tennis racket”. However, the input provided to EliGen is incorrect, which is featured two disjointed positions. In such scenarios, EliGen prioritizes the coherence of entities within the image, adaptively fine-tuning their actual positions, such as adjusting the position of the person’s hand. Though this adaption may result in the model underperforming on quantitative metrics related to spatial accuracy, it substantially improves the overall image quality.
@@ E.4 Pure Text-to-Image Generation Capability
Figure 16: Generalization tests with incorrect spatial relationships between entities. In the provided examples, the input positional relationships between entities are inaccurate. For instance, the entities ”person” and ”tennis racket” should intersect rather than being separated to reflect the relationship of a person holding a tennis racket. Under such cases, EliGen prioritizes ensuring the coherence of entities in the image, thereby adaptively fine-tuning the actual positions of the entities, such as adjusting the position of the person’s hand.

### v3

@@ Abstract
Recent advancements in diffusion models have significantly advanced text-to-image generation, yet global text prompts alone remain insufficient for achieving fine-grained control over individual entities within an image. To address this limitation, we present EliGen, a novel framework for Entity-level controlled image Generation. Firstly, we put forward regional attention, a mechanism for diffusion transformers that requires no additional parameters, seamlessly integrating entity prompts and arbitrary-shaped spatial masks. By contributing a high-quality dataset with fine-grained spatial and semantic entity-level annotations, we train EliGen to achieve robust and accurate entity-level manipulation, surpassing existing methods in both spatial precision and image quality. Additionally, we propose an inpainting fusion pipeline, extending EliGen’s capabilities to multi-entity image inpainting tasks. We further demonstrate EliGen’s flexibility by integrating it with other open-source models such as IP-Adapter, In-Context LoRA and MLLM, unlocking new creative possibilities. The source code, model, and dataset are published at the Project Page.
@@ 1 Introduction
Recently, diffusion models Rombach et al. (2022a); Esser et al. (2024); Labs (2024) have significantly advanced text-to-image generation, enabling the creation of high-quality images through textual descriptions. However, a single prompt is insufficient for precise image design, as natural language cannot describe image content with pixel-level precision, and image models often struggle to accurately model spatial relationships and object quantities Podell et al. (2024). This pitfall significantly limits their broader applicability. Consequently, integrating diverse conditions into text-to-image models has become a central research focus. For example, recent studies like ControlNet Zhang et al. (2023) and IP-Adapter Ye et al. (2023) have introduced conditions such as depth maps, normal maps and reference images to control the layout and style of generated images. Nevertheless, these conditions operate at a global image level, rather than targeting specific entities within the image. As a result, they lack the capability to provide fine-grained and entity-level control. This limitation impedes the generation of complex, highly customized images, motivating us to investigate precise controlling approaches for entity-level control.
@@ 2 Related Work
Text-to-image diffusion models Ho et al. (2020); Song et al. (2021); Li et al. (2024) have catalyzed the evolution of generative AI by generating high-quality images from textual descriptions. Latent Diffusion Models (LDMs) Rombach et al. (2022a); Podell et al. (2024) have emerged as a pivotal innovation, operating in a compressed latent space rather than directly in pixel space, which significantly reduces computational costs. These models primarily incorporate text conditions through the cross attention module within the UNet architecture Ronneberger et al. (2015). The Diffusion Transformer (DiT) Peebles and Xie (2022) represents the first diffusion model to utilize a transformer, integrating conditions into image latents via adaptive LayerNorm. SD3 Esser et al. (2024) and FLUX Labs (2024) are the state-of-the-art models that adopt the DiT, enhancing image generation by concatenating text and latent image modalities, thereby fully leveraging the semantic information of text to produce images of higher quality and finer detail. Despite the impressive capabilities of these models, generating images solely from text remains constrained. Recent studies have been exploring methods to introduce additional conditions Zhang et al. (2023); Mou et al. (2023); Ye et al. (2023), with entity-level control being one such sophisticated condition.
@@
To equip foundational models with fine-grained control capabilities, researchers originally explored train-free methods such as multi-diffusion Bar-Tal et al. (2023) and cross attention guidance Ma et al. (2024). However, these approaches often produce unstable results and encounter entity omission issues. To tackle these challenges, researchers introduced additional modules to encode spatial information for each entity using bounding boxes and conducted task-specific training. GLIGEN Li et al. (2023) is the first entity control model to undergo training, integrating an additional gated attention module and encoding bounding boxes through Fourier embedding. Subsequently, MIGC Zhou et al. (2024) introduces the shading aggregation controller and maps spatial information onto attention maps to guide cross attention processes. InstanceDiffusion Wang et al. (2024b) further categorizes spatial information from coarse to fine and designs unique tokenizers for each category. While these studies focus on designing complex structures to encode spatial information, this paper shifts the focus to designing and training regional attention mechanism without extra parameters and incorporating spatial information by controlling the regions where entity semantics exert influence.
@@ 3.1 Preliminaries
Entity-level controlled image generation aims at generating images based on both global and entity-level conditions. The diffusion model should ensure not only the accuracy of attributes and positions of each entity but also the overall harmony and quality of the image. The conditional inputs include a global text prompt $p$ and $L$ entity-level conditions. Each entity $e_{i}$ is composed of a local text prompt $p_{i}$ and spatial localization $l_{i}$, denoted as:
@@ 3.2 Regional Attention
Here, we construct the binary mask $\mathrm{M}$ to specify the regions where each token should attend during attention, as shown in Figure 3c, which primarily includes four categories of regions. For consistency, we treat the global prompt as a global “entity”, with its entity mask being 1 for the entire image.
@@ 3.3 Dataset with Entity Annotation
For the entity control task, each training sample should consist of an image, a global prompt, and multiple entity annotations. Regarding the image source, prior works typically use open-source datasets like Laion LAION (2022) for images, but we observe that Laion’s image distribution differs significantly from images generated by base model, leading to stylistic inconsistencies. Using Laion for training would distort the model distribution and degrade image quality (see the Supplementary Material for ablations), as the model would need to fit this mismatch. Therefore, we choose to directly use images generated by our baseline model, FLUX.1-dev, to avoid such distribution discrepancies.
The dataset generation pipeline and statistics are illustrated in the Supplementary Material. Starting with DiffusionDB Wang et al. (2022), we randomly select a caption and use FLUX.1-dev to generate the source image. Then, we employ Multimodal Large Language Model (MLLM) Wang et al. (2024a) to re-caption the image, forming the global prompt. Unlike previous methods relying on Grounding DINO Liu et al. (2023), we directly use Qwen2-VL 72B, which possesses the most powerful grounding capabilities across all MLLMs, to randomly select entities in the image and annotate their local prompts and localization (via rectangular masks). This ensures consistency between local and global prompts. Since our regional attention employs soft-coded spatial localization, training with rectangular masks is sufficient to enable the model to generalize to arbitrary-shaped masks.
@@ 4.1 Entity-Level Controlled Generation
In this scenario, entities are localized using bounding boxes. Our model converts bounding boxes to masks for input, while others utilize bounding box coordinates. The results are shown in Figure 4, where our model demonstrates superior performance in image quality, and entity accuracy. In the first two rows, simple tasks with distinct entities are handled by most methods except Multi-Diffusion. The last two rows highlight complex layouts where entities are intricately coupled. While other models struggle to maintain accuracy and coherence, our model consistently delivers high-quality and coherent images.
@@ 5.1 Experimental Setup
Following previous methods, we perform quantitative evaluations on the filtered COCO Lin et al. (2014); Zheng et al. (2023) 2017 validation set, comprising 2830 prompt-image pairs. Previous GLIGEN Li et al. (2023) utilized the mean average precision for evaluation, a metric that computes the area under the precision-recall curve. However, we observe its critical limitation: it treats background entities in generated images as negative samples, resulting in substantial inaccuracies. As a result, we decide not to use this metric. Instead, to thoroughly evaluate the models’ capabilities on the COCO benchmark, we assess their performance across the following dimensions.
@@ 5.2 Comparison on COCO Benchmark
The evaluation results of the models on the COCO benchmark are presented in Table 1. For Entity Success Rate, we assessed all methods using Qwen2-VL Wang et al. (2024a), InternVL2 Chen et al. (2024), and CogVLM2 Hong et al. (2024). The results demonstrate that EliGen consistently outperforms other models, highlighting its robustness and high quality. In terms of mIoU, our model performs comparably to Instance-Diffusion and substantially better than other methods. The slightly lower mIoU of EliGen compared to Instance-Diffusion is primarily due to its soft-coding scheme for positional information. This scheme allows EliGen to automatically expand certain regions to ensure image coherence, as further illustrated in the Supplementary Material. In terms of both CLIP Score and Aesthetic Score, EliGen significantly outperforms other models, demonstrating substantial superiority in image quality.
@@ 6 Conclusion
To achieve image generation with entity-level controls, we introduce regional attention into the DiT architecture to effectively integrate entity prompts and their locations. Furthermore, we construct a dataset with fine-grained entity information and finetune the regional attention mechanism, developing the EliGen model. EliGen demonstrates high generalization in entity-level controlled image generation, delivering superior image quality and entity accuracy compared to existing models. We further propose an inpainting fusion pipeline, extending EliGen’s functionality to image inpainting tasks. Moreover, by integrating EliGen with open-source models, we demonstrate its potential for creative applications, broadening its scope and versatility.
@@ Appendix B Inpainting Fusion Algorithm
Figure 11: Dialogue-based image design and editing with MLLM and our EliGen. Users can utilize MLLM to design the layout of the image that they wish to generate or input the image that they want to edit into MLLM. After MLLM generates the caption and entity information of the image, EliGen can produce the target image.
@@ Appendix C Potential Creative Extension
Figure 12: Training qualitative comparison between our constructed dataset and Laion dataset. The image pair is generated from the same noise, and the global prompt is “A person standing by the river”. The comparative results show that the images generated by the model trained on our proposed dataset significantly outperform those from the Laion dataset in quality.
@@ C.2 Entity Transfer with In-Context LoRA
In-Context LoRA Huang et al. [2024a] is capable of performing few-shot image composition tasks conditioned on either global prompts or images, and can be applied in various scenarios such as portrait photography, product design, font design, and visual identity design. Our EliGen model further enhances these scenarios by incorporating entity-level control conditions, enabling more personalized image generation. Taking visual identity design as an example, we have successfully accomplished the entity transfer task by introducing EliGen, as illustrated in Figure 10. Given the entity defined by the source image, we can transfer it to any desired location in the target image through EliGen’s positional control capability, thereby achieving the intended visual identity.
@@ C.3 Dialogue-based Image Design and Editing with MLLM
The MLLM is capable of interpreting user-provided images and textual content to engage in dialogue. As a result, we can achieve dialogue-based image design and editing by leveraging the integration of MLLM and EliGen. Specifically, grounding-enabled MLLMs such as Qwen2-VL Wang et al. [2024a] can be utilized to determine entity locations based on textual descriptions, enabling the generation of images solely through text using EliGen. Furthermore, the MLLM can be employed to comprehend images and design entity information in accordance with specific image editing requirements, thereby facilitating precise image editing in conjunction with EliGen. The qualitative results are shown in Figure 11.
@@ D.1 Training Qualitative Comparison for Dataset Image Source
The primary objective of training the EliGen model is to activate its capability for layout control, which necessitates the construction of a dataset annotated with entities. Moreover, the image distribution within the dataset should closely align with the output distribution of the model itself. If the distribution gap is substantial, a significant portion of the training loss would be attributed to this discrepancy. This would significantly impede the model’s convergence and degrade the quality of the generated images. Consequently, we opt to utilize the Flux-generated images for dataset creation, rather than directly employing images from the open-source Laion dataset LAION [2022].
We conduct training experiments using both our custom-created dataset and the Laion dataset, where the annotations for the Laion dataset are derived from the same pipeline described in Section A.1. As illustrated in Figure 12, under identical prompts and initial noise conditions, the model trained on our dataset demonstrates significantly superior image quality compared to the model trained on the Laion dataset. Compared to Flux-generated images, the latter exhibits significant visual clutter, inferior detail quality, and a notable deviation from the expected distribution.
@@ E.2 Continuously Varying Entity Positions
In Figure 15, we evaluate the control efficacy of EliGen using identical input noise and the same entity. The position of the “person” entity shifts progressively from left to right, with EliGen maintaining consistent control throughout the process. To preserve the identity of the person during generation, we integrate the IP-Adapter Team [2024] method. This demonstrates EliGen’s potential to generate coherent image sequences, such as those involving continuous movement of entities through positional control or complex actions orchestrated via coordinated control of multiple entities.
@@ E.3 Unreasonable Positional Inputs
incorrect spatial relationships between entities refer to scenarios where the relative positions of entities are inaccurate, such as being too far apart or too close together. We evaluate two cases of this inputs, as depicted in Figure 16. In both examples, the correct positioning of the “person” entity necessitates an overlap with another entity to accurately convey the action being performed. For example, in the prompt “A person is playing tennis”, the “person” should ideally be holding a “tennis racket”. However, the input provided to EliGen is incorrect, which is featured two disjointed positions. In such scenarios, EliGen prioritizes the coherence of entities within the image, adaptively fine-tuning their actual positions, such as adjusting the position of the person’s hand. Though this adaption may result in the model underperforming on quantitative metrics related to spatial accuracy, it substantially improves the overall image quality.
@@ E.4 Pure Text-to-Image Generation Capability
Figure 16: Generalization tests with incorrect spatial relationships between entities. In the provided examples, the input positional relationships between entities are inaccurate. For instance, the entities “person” and “tennis racket” should intersect rather than being separated to reflect the relationship of a person holding a tennis racket. Under such cases, EliGen prioritizes ensuring the coherence of entities in the image, thereby adaptively fine-tuning the actual positions of the entities, such as adjusting the position of the person’s hand.